According to a Pew Research Center study, in May 2011, 30% of all American adults had a smart phone (one which the user can use to read email and surf the Internet). A communications professor at a university believes this percentage is higher among community college students. She selects 441 community college students at random and finds that 140 of them have a smart phone. Then in testing the hypotheses:

H0: p = 0.3 versus

Ha: p > 0.3,

what is the test statistic?

z =  . (Please round your answer to two decimal places.)
Answer #1

This is the right tailed test .

The null and alternative hypothesis is

H0 : p =0.30

Ha : p >0.30

\hat p = x / n = 140/441=0.32

P0 = 0.30

1 - P0 = 0.70

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.32-0.30/ [\sqrt(0.30*0.70) / 441]

= 0.92